# The Most Common Habitable Planets III – Modeling Temperature Forcing and Surface Conditions on Rocky Exoplanets and Exomoons

### v1

@@ Abstract
Small rocky planets, as well as larger planets that suffered extensive volatile loss, tend to be drier and have thinner atmospheres as compared to Earth. Such planets probably outnumber worlds better endowed with volatiles, being the most common habitable planets. For the subgroup of fast rotators following eccentric orbits, atmospheres suffer radiative forcing and their heat capacity provides a method for gauging atmospheric thickness and surface conditions. We further explore the model presented in a previous paper and apply it to real and hypothetical exoplanets in the habitable zone of various classes of stars, simulating atmospheric and orbital characteristics. For planetary eccentricities e $\sim$0.3, the forcing-induced hypothetical temperature variation would reach $\sim 80$ K for airless planets and $\sim 10$ K for planets with substantial atmospheres. For Kepler-186 f and Kepler-442 b, assuming e $\sim$0.1, temperature variations can reach $\sim 24$ K. We also consider habitable exomoons in circular orbits around gas giants within the habitable zone, which suffer radiative forcing due to their epicyclic motion. We study several combinations of parameters for the characterization of planets (mass, eccentricity and semi-major axis) and exomoons (mass, orbital radius, albedo and atmospheric characteristics) for different stellar types. For e $\sim$0.3, exomoon temperature varies up to $\sim 90$ K, while for $\sim$0.6 variations can reach $\sim 200$ K. Such exomoons may plausibly retain their volatiles by continued volcanic activity fueled by tidal dissipation. Although currently undetectable, such effects might be within reach of future ELT-class telescopes and space missions with mid-infrared and coronagraphic capabilities.
@@ 1 Introduction
Since the discovery of the first exoplanets in the 90s 83; 51, more than 5600 other planets outside the Solar System were confirmed through different observation techniques (http://exoplanet.eu/), mainly due to the Kepler space missions 4. Many of them are part of multiple systems, and current and future missions such as the Transiting Exoplanet Survey Satellite (TESS), 28; 46; 21; 39; 68; 7; 79; 54 and the Planetary Transits and Oscillations of stars (PLATO) 66 are expected to discover many more.
Many such planets lie within the habitable zone (hereafter HZ) of long lived stars 78; 11; 17. Systems such as TRAPPIST-1, with seven Earth-sized planets orbiting an M dwarf star 22, three of which believed to be inside the HZ 27, naturally spark our curiosity on the possibility of such systems harboring life. The existence of life, based on our experience on Earth, depends on the existence of water in the liquid state. So far, based on the forms of life we know of, it seems that a rocky exoplanet with stable suitable temperatures and an atmosphere, compatible with liquid water at the surface, would be the most interesting candidate to look for biosignatures.
@@
As shown in Pinotti 59, the presence of an atmosphere contributes to buffer brightness temperature variations due to eccentric orbits, in planets which rotate fast enough so that temperature variations between night and day time can be disregarded. The fraction of the atmosphere which responds radiatively will have a thermal capacitance, and its buffering property will eventually be sensed at the surface level, whose temperature takes into account the greenhouse effect. The brightness temperature will coincide with surface temperature only for the case of no atmosphere. This attenuation and its associated phase lag in relation to the stellar flux profile could, in principle, be detectable through infrared photometric time series, allowing an estimate of atmospheric thickness, as well as, possibly, other atmospheric properties such as gross composition. Therefore, the modeling of such temperature variations can be effective in characterizing the surface conditions of rocky exoplanets when future instrumentation enables the measurement of their mid-infrared emission, through coronagraphic or interferometric techniques, for example ELT/METIS 5.
@@
The aim of this paper is two-fold: a wider exploration of the model presented in Pinotti 59 and its application to various sets of real and hypothetical systems. We study the stability of the brightness temperature on different rocky exoplanets, exploring a wide parameter space in both stellar and planetary properties, and discussing on how a periodic temperature profile measurement could be used to indicate the presence of an atmosphere, and to infer some of its main characteristics as well as help establish if a rocky planet is able to maintain liquid water at its surface.
@@
We also study brightness temperature variations for hypothetical rocky exomoons orbiting Jupiter sized exoplanets inside what would be the HZ for a rocky planet. Planetary migration is a common phenomenon, deduced from their frequently very small semi-major axis 41; 80; 50; 30. Volatile-rich moons are expected to lose most if not all of their volatile content as they migrate to the HZ, since their low mass is not capable of holding on to substantial atmospheres. Another contributing factor is the extended pre-main sequence and early main-sequence phase of higher luminosity that affects low mass stars before they settle onto the main sequence. This extended phase would keep exoplanets and their moons within conditions of rapid water loss for many times $\sim$108 years and may be relevant for stars as massive as K5 63. For exoplanets about the size of the Earth, complete water loss ensues (unless the planet is endowed with a very large initial water inventory) and it is to be expected that exomoons of giant exoplanets, close to or within the HZ, will suffer a similar fate 76; 45. However, for exomoons tidal dissipation as they pursue their fast orbits, maybe tied to relationships with other moons as well, might keep the moon’s internal engine going on long and strongly enough to enable protracted volcanism, atmosphere replenishment and crustal rejuvenation. It is thus to be expected that massive, rocky or partly rocky exomoons might retain a measurable atmosphere, even after migrating inward from their original place of formation at or beyond the snowline. The model of Pinotti 59 has been upgraded in order to take into account the convolution of the orbit of an exoplanet and its exomoon, and it proposes that the determination of their brightness temperatures could be instrumental in characterizing the exomoons’ surface conditions and habitability for a wide range of realistic conditions. Although not yet detected, exomoons are expected to be abundant and are already a popular debate subject among the recent literature 70; 13; 89; 20, including the recent suggestion of an “exomoon” orbiting a nearby brown dwarf 19. Large rocky exomoons with an atmosphere could be as likely to harbor life as a similar sized exoplanet 26; 23. In fact, for the case of compact planetary orbits, where tidal locking between planet and star is thought to be a possible cause of significant temperature variations on the planet’s surface 24; 18, should not be a problem for the stabilization of the climate on an exomoon, provided that the orbital periods remain not too far from a few terrestrial days, so they can be considered fast rotators.
@@ 2 Temperature variation model
The following mathematical development is based on Pinotti 59, where its details can be appreciated. It considers the atmosphere of a planet or a moon as divided between an outer part, whose main heat exchange mechanism is radiative, and an inner part, dominated by convection. The thermal radiation of the outer part is detectable through its brightness temperature, $T$. For a planet or a moon subjected to temporal variation of the stellar radiation flux, $F_{\star}$, the brightness temperature responds according to
@@
Several limitations of the model must be kept in mind. Firstly, Eq. (1) assumes that the planet or moon is a fast rotator, that is, the radiative timescale of the atmosphere is larger than the rotation period. Secondly, the model treats radiation bolometrically, in a simplified way, with no regard to the details of the radiative transfer and its spectral distribution. Thirdly, it further assumes that the planet or moon is volatile deficient so that energy transfer in absorption, evaporation and condensation does not have to be taken into account, and the albedo, which relates to a cloud coverage, which in turn depends on volatile content in the surface of the planet or moon, does not fluctuate considerably along an eccentric orbit. The model also works if a planet rich in volatiles has an average surface temperature high enough so that the volatiles are mostly in the gaseous phase. While the problem of the origin of the Earth’s oceans remains unresolved, with arguments for both wet-endogenous and exogenous sources 91; 8; 31, many researchers indicate that prolonged exposure to XUV radiation and particle winds, especially for planets around low-mass stars, may promote substantial volatile loss 67; 1; 3 from rocky planets.
Furthermore, rocky planets inside the HZ of M stars may also be intrinsically water deficient from the point of view of wet-endogenous source (43, but see, for example, 2 for alternative scenarios). More severe volatile loss is expected for low mass planets, due to their lower gravitational fields, and also due to their weaker magnetosphere 90. Therefore, Eq. 1 applies best to planets or moons smaller than the Earth, for which volatile scarcity and thin atmospheres should be a common condition. Even planets as massive as the Earth or even larger may frequently incur in severe atmospheric loss and thus be describable by our model. A more recent discussion on volatile loss and water origin in rocky planets can be found in Pinotti & Porto de Mello 60.
Studies of planetary atmospheres subjected to radiative forcing from their parent stars, and with sophisticated tools like Global Circulation Models 35; 56; 82 are usually focused on Earth-like planets, rich in volatiles, having, therefore, more complex energy fluxes, including water evaporation and condensation. Besides, the focus of those studies is on habitability conditions only. Our work, on the other hand, is devoted mainly to an observational method for estimating the atmospheric thickness of planets or moons with low volatile content, which may well be the majority in the Universe.
@@
where $\bar{T}=T-T(a)$, $K$ is the system gain ($K=[4\,\sigma\,\varepsilon(T(a)^{3})]^{-1})$, $H$ is the amplitude of the forcing, $\tau$ is the time constant of the linearized form of Eq. 1 (see 59), and $\phi$ is the phase lag ($\phi=-{\rm tan}^{-1}(\omega\,\tau$)). We see then that the amplitude of $\bar{T}$ is affected by the term $(\omega^{2}\,\tau^{2}+1)^{-0.5}$, that is, the attenuation or damping of the amplitude of $\bar{T}$ is related to $\omega$, which in turn depends strongly on the value of $a$ (see Eq. 4). That is why the damping will be more evident for planets in the HZ of later types stars. The attenuation is also important in terms of habitability, and this relation is discussed in Pinotti 59.
@@ 3 A hypothetical Earth: a case study
In order to further understand our model’s behavior and its dependence on each free parameter, we applied it to hypothetical Earth analogs. For these hypothetical Earths, we solved Eq. (1) while varying the value of each parameter individually and then studied their influence on the temperature profile through four indicators: attenuation, phase lag (both already studied in Pinotti 59), asymmetry, and narrowing, proposed in the current paper.
@@
Aiming to study how the orbital eccentricity influences the behavior of a planet’s brightness temperature, we solved Eq. (1) for four distinct eccentricity values: $e=$ 0.0167, 0.1, 0.2, and 0.3. As exhibited in Fig. 2, for highly eccentric orbit, our hypothetical Earth based planet can reach a temperature variation of 80 K throughout the year. Since these planets have no atmosphere, there will be no phase lag or attenuation. One can see that the eccentricity does not cause an asymmetry in the curve, but does cause a narrowing, that increases proportionately: the more elliptical the orbit, the sharper the temperature increase near periastron, a well-known effect that drives Martian global dust storms since southern summer takes place close to maximum proximity to the Sun 72. The remaining parameters were studied only with the two extreme values of eccentricity (0.0167 and 0.3).
Figure 2: Temperature as a function of time for four hypothetical planets without an atmosphere ($\chi=0$), $A=0.3$, $\epsilon=0.9$ and distinct values of eccentricity: 0.0167 (brown), 0.1 blue), 0.2 (orange), and 0.3 (cyan).
@@
In order to study the influence of the atmospheric heat capacity and column density, we included an atmosphere to our model, with template values of cp = 32 J mol-1 K-1 and $\chi$ = 3 ×$\times 10^{4}$ mol m-2. The atmospheric column density value is the same as Earth’s, obtained by assuming the radiative part of the atmosphere to start above a pressure value of 0.1 bar 69 and exponential profiles for density and pressure, with scale height $h=8.5$ km for both (taken from NASA’s “fact sheet" webpage [^1] https://nssdc.gsfc.nasa.gov/planetary/planetfact.html). We decided to use this value of heat capacity based on Pinotti 59 that considers the sensitivity of heat capacity for relevant gases of rocky planets’ atmospheres (CH4, O2, N2, NH3, H2O, CO2 and H2S), which vary between 29.1 J mol-1 K-1 and 37.3 J mol-1 K-1 for 300 K. We checked this is valid as well for H2 rich atmospheres, as might be the case in planets with intense volcanic activity 58; 64, since H2 has cp = 28.8 J mol-1 K-1 at 300 K. The heat capacity and the column density are two important variables that can influence the degree of damping of the upper atmosphere’s temperature.
@@
Figure 3: Brightness temperature as a function of time for hypothetical planets with $A=0.3$ and $\varepsilon=0.9$. The dotted brown curves show the cases with no atmosphere, for comparison, and the solid lines show the cases for an atmosphere with $\chi=3\times 10^{4}$ mol m-2 for the upper panel and $\chi=3\times 10^{5}$ mol m-2 for the lower one. The planets have two distinct values of eccentricity: 0.0167 (smallest amplitude curves) and 0.3 (curves with highest amplitude). For the cases with an atmosphere, the heat capacity is either 30 J mol-1 K-1 (in blue) or 90 J mol-1 K-1 (in orange).
@@
As already verified, the value of atmospheric column density $\chi$ can have a big impact on the temperature profile. In order to further study the influence of $\chi$, we solved Eq. (1) using three different values of $\chi$: the first based on Mars’ atmosphere ($\chi=5\times 10^{3}$ mol m-2), the second based on Earth’s ($\chi$ = $3\times 10^{4}$ mol m-2), and the third one based on Titan’s ($\chi=5\times 10^{6}$ mol m-2). The values of $\chi$ for Mars and Titan were obtained assuming exponential profiles for density and pressure, with the same value of $h$ for both, as was done for Earth. We considered the thin Martian atmosphere to be fully radiative, with $h=11.1$ km (taken from NASA’s “fact sheet" webpage). For Titan, we considered the transition between a convective and radiative atmosphere to occur when pressure reaches 1.3 bar. This value is much higher than the 0.1 bar adopted for Earth (which can also be assumed for other planets) because shortwave absorption in Titan’s upper hazy troposphere causes stability against convection 69. Using values of $h$ between 15 km and 50 km 29, we obtained the average value of $\chi=5\times 10^{6}$ mol m-2 for Titan. In Fig. 4 we can see the temperature profiles for planets with different values of atmospheric $\chi$. We chose to display only the cases for the atmospheres of Earth and Titan, since the results for Mars’ atmosphere do not produce any significant changes from the Earth’s. Actually, we have verified that, keeping all the other parameters unchanged, significant changes from the temperature profile obtained for $\chi=3\times 10^{4}$ mol m-2 (the value for Earth’s atmosphere) appear only for atmospheres with $\chi\gtrsim 3\times 10^{5}$ mol m-2. Fig. 4 shows the temperature profiles for the two eccentricity cases we have selected compared to the same planet without an atmosphere. All the curves were analyzed with the four indicators. The data shows that the presence of an atmosphere not only strongly attenuates the temperature variation but also causes a phase lag. A thicker atmosphere also increases the asymmetry, stressing the elliptical nature of the orbit, and decreases the narrowing, which causes the curve to be smoother.
@@
Figure 4: Brightness temperature as a function of time for two hypothetical planets with an atmosphere (solid curves) compared to the corresponding cases without an atmosphere (dotted, brown curves). The planets have two distinct values of eccentricity: 0.0167 (smallest amplitude curves) and 0.3 (curves with highest amplitude) and fixed $\varepsilon=0.9$, $A=0.3$ and cp = 32 J mol-1 K-1. For the cases with an atmosphere, the column density is either $\chi=3\times 10^{4}$ mol m-2 (in blue) or $\chi=5\times 10^{6}$ mol m-2 (in orange).

### v2

@@ Abstract
Small rocky planets, as well as larger planets that suffered extensive volatile loss, tend to be drier and have thinner atmospheres as compared to Earth. Such planets probably outnumber worlds better endowed with volatiles, being the most common habitable planets. For the subgroup of fast rotators following eccentric orbits, atmospheres suffer radiative forcing and their heat capacity provides a method for gauging atmospheric thickness and surface conditions. We further explore the model presented in a previous paper and apply it to real and hypothetical exoplanets in the habitable zone of various classes of stars, simulating atmospheric and orbital characteristics. For planetary eccentricities e $\sim$0.3, the forcing-induced hypothetical temperature variation would reach $\sim 80$ K for airless planets and $\sim 10$ K for planets with substantial atmospheres. For Kepler-186 f and Kepler-442 b, assuming e $\sim$0.1, temperature variations can reach $\sim 24$ K. We also consider habitable exomoons in circular orbits around gas giants within the habitable zone, which suffer radiative forcing due to their epicyclic motion. We study several combinations of parameters for the characterization of planets (mass, eccentricity and semi-major axis) and exomoons (mass, orbital radius, albedo and atmospheric characteristics) for different stellar types. For e $\sim$0.3, exomoon temperature varies up to $\sim 90$ K, while for $\sim$0.6 variations can reach $\sim 200$ K. Such exomoons may plausibly retain their volatiles by continued volcanic activity fueled by tidal dissipation. Although currently undetectable, such effects might be within reach of future Extremely Large Telescope-class telescopes and space missions with mid-infrared and coronagraphic capabilities.
@@ 1 Introduction
Since the discovery of the first exoplanets in the 90s 81; 49, more than 5600 other planets outside the Solar System were confirmed through different observation techniques (http://exoplanet.eu/), mainly due to the Kepler space missions 4. Many of them are part of multiple systems, and current and future missions such as the Transiting Exoplanet Survey Satellite (TESS), 27; 44; 20; 37; 66; 7; 77; 52 and the Planetary Transits and Oscillations of stars (PLATO) 64 are expected to discover many more.
Many such planets lie within the habitable zone (hereafter HZ) of long lived stars 76; 11; 16. Systems such as TRAPPIST-1, with seven Earth-sized planets orbiting an M dwarf star 21, three of which believed to be inside the HZ 26, naturally spark our curiosity on the possibility of such systems harboring life. The existence of life, based on our experience on Earth, depends on the existence of water in the liquid state. So far, based on the forms of life we know of, it seems that a rocky exoplanet with stable suitable temperatures and an atmosphere, compatible with liquid water at the surface, would be the most interesting candidate to look for biosignatures.
@@
As shown in Pinotti 57, the presence of an atmosphere contributes to buffer brightness temperature variations due to eccentric orbits, in planets which rotate fast enough so that temperature variations between night and day time can be disregarded. The fraction of the atmosphere which responds radiatively will have a thermal capacitance, and its buffering property will eventually be sensed at the surface level, whose temperature takes into account the greenhouse effect. The brightness temperature will coincide with surface temperature only for the case of no atmosphere. This attenuation and its associated phase lag in relation to the stellar flux profile could, in principle, be detectable through infrared photometric time series, allowing an estimate of atmospheric thickness, as well as, possibly, other atmospheric properties such as gross composition. Therefore, the modeling of such temperature variations can be effective in characterizing the surface conditions of rocky exoplanets when future instrumentation enables the measurement of their mid-infrared emission, through coronagraphic or interferometric techniques, for example Extremely Large Telescope (ELT) / Mid-infrared ELT Imager and Spectrograph (METIS) 5.
@@
The aim of this paper is two-fold: a wider exploration of the model presented in Pinotti 57 and its application to various sets of real and hypothetical systems. We study the stability of the brightness temperature on different rocky exoplanets, exploring a wide parameter space in both stellar and planetary properties, and discussing on how a periodic temperature profile measurement could be used to indicate the presence of an atmosphere, and to infer some of its main characteristics as well as help establish if a rocky planet is able to maintain liquid water at its surface.
@@
We also study brightness temperature variations for hypothetical rocky exomoons orbiting Jupiter sized exoplanets inside what would be the HZ for a rocky planet. Planetary migration is a common phenomenon, deduced from their frequently very small semi-major axis 39; 78; 48; 29. Volatile-rich moons are expected to lose most if not all of their volatile content as they migrate to the HZ, since their low mass is not capable of holding on to substantial atmospheres. Another contributing factor is the extended pre-main sequence and early main-sequence phase of higher luminosity that affects low mass stars before they settle onto the main sequence. This extended phase would keep exoplanets and their moons within conditions of rapid water loss for many times $\sim$108 years and may be relevant for stars as massive as K5 61. For exoplanets about the size of the Earth, complete water loss ensues (unless the planet is endowed with a very large initial water inventory) and it is to be expected that exomoons of giant exoplanets, close to or within the HZ, will suffer a similar fate 74; 43. However, for exomoons tidal dissipation as they pursue their fast orbits, maybe tied to relationships with other moons as well, might keep the moon’s internal engine going on long and strongly enough to enable protracted volcanism, atmosphere replenishment and crustal rejuvenation. It is thus to be expected that massive, rocky or partly rocky exomoons might retain a measurable atmosphere, even after migrating inward from their original place of formation at or beyond the snowline. The model of Pinotti 57 has been upgraded in order to take into account the convolution of the orbit of an exoplanet and its exomoon, and it proposes that the determination of their brightness temperatures could be instrumental in characterizing the exomoons’ surface conditions and habitability for a wide range of realistic conditions. Although not yet detected, exomoons are expected to be abundant and are already a popular debate subject among the recent literature 68; 13; 87; 19, including the recent suggestion of an “exomoon” orbiting a nearby brown dwarf 18. Large rocky exomoons with an atmosphere could be as likely to harbor life as a similar sized exoplanet 25; 22. In fact, for the case of compact planetary orbits, where tidal locking between planet and star is thought to be a possible cause of significant temperature variations on the planet’s surface 23; 17, should not be a problem for the stabilization of the climate on an exomoon, provided that the orbital periods remain not too far from a few terrestrial days, so they can be considered fast rotators.
@@ 2 Temperature variation model
The following mathematical development is based on Pinotti 57, where its details can be appreciated. It considers the atmosphere of a planet or a moon as divided between an outer part, whose main heat exchange mechanism is radiative, and an inner part, dominated by convection. The thermal radiation of the outer part is detectable through its brightness temperature, $T$. For a planet or a moon subjected to temporal variation of the stellar radiation flux, $F_{\star}$, the brightness temperature responds according to
@@
Several limitations of the model must be kept in mind. Firstly, Eq. (1) assumes that the planet or moon is a fast rotator, that is, the radiative timescale of the atmosphere is larger than the rotation period. Secondly, the model treats radiation bolometrically, in a simplified way, with no regard to the details of the radiative transfer and its spectral distribution. Thirdly, it further assumes that the planet or moon is volatile deficient so that energy transfer in absorption, evaporation and condensation does not have to be taken into account, and the albedo, which relates to a cloud coverage, which in turn depends on volatile content in the surface of the planet or moon, does not fluctuate considerably along an eccentric orbit. The model also works if a planet rich in volatiles has an average surface temperature high enough so that the volatiles are mostly in the gaseous phase. While the problem of the origin of the Earth’s oceans remains unresolved, with arguments for both wet-endogenous and exogenous sources 89; 8; 30, many researchers indicate that prolonged exposure to XUV radiation and particle winds, especially for planets around low-mass stars, may promote substantial volatile loss 65; 1; 3 from rocky planets.
Furthermore, rocky planets inside the HZ of M stars may also be intrinsically water deficient from the point of view of wet-endogenous source (41, but see, for example, 2 for alternative scenarios). More severe volatile loss is expected for low mass planets, due to their lower gravitational fields, and also due to their weaker magnetosphere 88. Therefore, Eq. 1 applies best to planets or moons smaller than the Earth, for which volatile scarcity and thin atmospheres should be a common condition. Even planets as massive as the Earth or even larger may frequently incur in severe atmospheric loss and thus be describable by our model. A more recent discussion on volatile loss and water origin in rocky planets can be found in Pinotti & Porto de Mello 58.
Studies of planetary atmospheres subjected to radiative forcing from their parent stars, and with sophisticated tools like Global Circulation Models 54; 80 are usually focused on Earth-like planets, rich in volatiles, having, therefore, more complex energy fluxes, including water evaporation and condensation. Besides, the focus of those studies is on habitability conditions only. Our work, on the other hand, is devoted mainly to an observational method for estimating the atmospheric thickness of planets or moons with low volatile content, which may well be the majority in the Universe.
@@
where $\bar{T}=T-T(a)$, $K$ is the system gain ($K=[4\,\sigma\,\varepsilon(T(a)^{3})]^{-1})$, $H$ is the amplitude of the forcing, $\tau$ is the time constant of the linearized form of Eq. 1 (see 57), and $\phi$ is the phase lag ($\phi=-{\rm tan}^{-1}(\omega\,\tau$)). We see then that the amplitude of $\bar{T}$ is affected by the term $(\omega^{2}\,\tau^{2}+1)^{-0.5}$, that is, the attenuation or damping of the amplitude of $\bar{T}$ is related to $\omega$, which in turn depends strongly on the value of $a$ (see Eq. 4). That is why the damping will be more evident for planets in the HZ of later types stars. The attenuation is also important in terms of habitability, and this relation is discussed in Pinotti 57.
@@ 3 A hypothetical Earth: a case study
In order to further understand our model’s behavior and its dependence on each free parameter, we applied it to hypothetical Earth analogs. For these hypothetical Earths, we solved Eq. (1) while varying the value of each parameter individually and then studied their influence on the temperature profile through four indicators: attenuation, phase lag (both already studied in Pinotti 57), asymmetry, and narrowing, proposed in the current paper.
@@
Aiming to study how the orbital eccentricity influences the behavior of a planet’s brightness temperature, we solved Eq. (1) for four distinct eccentricity values: $e=$ 0.0167, 0.1, 0.2, and 0.3. As exhibited in Fig. 2, for highly eccentric orbit, our hypothetical Earth based planet can reach a temperature variation of 80 K throughout the year. Since these planets have no atmosphere, there will be no phase lag or attenuation. One can see that the eccentricity does not cause an asymmetry in the curve, but does cause a narrowing, that increases proportionately: the more elliptical the orbit, the sharper the temperature increase near periastron, a well-known effect that drives Martian global dust storms since southern summer takes place close to maximum proximity to the Sun 70. The remaining parameters were studied only with the two extreme values of eccentricity (0.0167 and 0.3).
Figure 2: Temperature as a function of time for four hypothetical planets without an atmosphere ($\chi=0$), $A=0.3$, $\epsilon=0.9$ and distinct values of eccentricity: 0.0167, 0.1, 0.2 and 0.3.
@@
In order to study the influence of the atmospheric heat capacity and column density, we included an atmosphere to our model, with template values of cp = 32 J mol-1 K-1 and $\chi$ = 3 ×$\times 10^{4}$ mol m-2. The atmospheric column density value is the same as Earth’s, obtained by assuming the radiative part of the atmosphere to start above a pressure value of 0.1 bar 67 and exponential profiles for density and pressure, with scale height $h=8.5$ km for both (taken from NASA’s “planetary fact sheet" webpage [^1] https://nssdc.gsfc.nasa.gov/planetary/planetfact.html). We decided to use this value of heat capacity based on Pinotti 57 that considers the sensitivity of heat capacity for relevant gases of rocky planets’ atmospheres (CH4, O2, N2, NH3, H2O, CO2 and H2S), which vary between 29.1 J mol-1 K-1 and 37.3 J mol-1 K-1 for 300 K. We checked this is valid as well for H2 rich atmospheres, as might be the case in planets with intense volcanic activity 56; 62, since H2 has cp = 28.8 J mol-1 K-1 at 300 K. The heat capacity and the column density are two important variables that can influence the degree of damping of the upper atmosphere’s temperature.
@@
Figure 3: Brightness temperature as a function of time for hypothetical planets with $A=0.3$ and $\varepsilon=0.9$. The dotted brown curves show the cases with no atmosphere, for comparison, and the solid lines show the cases for an atmosphere with $\chi=3\times 10^{4}$ mol m-2 for the upper panel and $\chi=3\times 10^{5}$ mol m-2 for the lower one. The planets have two distinct values of eccentricity: 0.0167 (smallest amplitude curves) and 0.3 (curves with highest amplitude). For the cases with an atmosphere, the heat capacity is either 30 J mol-1 K-1 or 90 J mol-1 K-1.
@@
As already verified, the value of atmospheric column density $\chi$ can have a big impact on the temperature profile. In order to further study the influence of $\chi$, we solved Eq. (1) using three different values of $\chi$: the first based on Mars’ atmosphere ($\chi=5\times 10^{3}$ mol m-2), the second based on Earth’s ($\chi$ = $3\times 10^{4}$ mol m-2), and the third one based on Titan’s ($\chi=5\times 10^{6}$ mol m-2). The values of $\chi$ for Mars and Titan were obtained assuming exponential profiles for density and pressure, with the same value of $h$ for both, as was done for Earth. We considered the thin Martian atmosphere to be fully radiative, with $h=11.1$ km (taken from NASA’s “planetary fact sheet" webpage). For Titan, we considered the transition between a convective and radiative atmosphere to occur when pressure reaches 1.3 bar. This value is much higher than the 0.1 bar adopted for Earth (which can also be assumed for other planets) because shortwave absorption in Titan’s upper hazy troposphere causes stability against convection 67. Using values of $h$ between 15 km and 50 km 28, we obtained the average value of $\chi=5\times 10^{6}$ mol m-2 for Titan. In Fig. 4 we can see the temperature profiles for planets with different values of atmospheric $\chi$. We chose to display only the cases for the atmospheres of Earth and Titan, since the results for Mars’ atmosphere do not produce any significant changes from the Earth’s. Actually, we have verified that, keeping all the other parameters unchanged, significant changes from the temperature profile obtained for $\chi=3\times 10^{4}$ mol m-2 (the value for Earth’s atmosphere) appear only for atmospheres with $\chi\gtrsim 3\times 10^{5}$ mol m-2. Fig. 4 shows the temperature profiles for the two eccentricity cases we have selected compared to the same planet without an atmosphere. All the curves were analyzed with the four indicators. The data shows that the presence of an atmosphere not only strongly attenuates the temperature variation but also causes a phase lag. A thicker atmosphere also increases the asymmetry, stressing the elliptical nature of the orbit, and decreases the narrowing, which causes the curve to be smoother.
@@
Figure 4: Brightness temperature as a function of time for two hypothetical planets with an atmosphere (solid curves) compared to the corresponding cases without an atmosphere (dotted, brown curves). The planets have two distinct values of eccentricity: 0.0167 (smallest amplitude curves) and 0.3 (curves with highest amplitude) and fixed $\varepsilon=0.9$, $A=0.3$ and cp = 32 J mol-1 K-1. For the cases with an atmosphere, the column density is either $\chi=3\times 10^{4}$ mol m-2 or $\chi=5\times 10^{6}$ mol m-2.